High-Low Method Quality Tools. Quality Tools Incorporated would like to estimate costs associated with its sales personnel. Salespeople are paid a salary plus commission. Commission rates vary among products and are based on sales dollars. The company reported the following monthly cost data related to sales personnel:

Reporting Period (Month) Total Costs Sales Amount
January $710,000 $13,800,000
February 695,000 13,600,000
March 765,000 15,100,000
April 650,000 12,000,000
May 775,000 15,500,000
June 750,000 14,700,000
July 715,000 14,500,000
August 680,000 13,100,000
September 830,000 16,500,000
October 815,000 16,000,000
November 800,000 15,600,000
December 690,000 13,200,000

Required:

  1. Use the four steps of the high-low method to estimate total fixed costs per month and the variable cost per sales dollar. State your results in the cost equation form Y = f + vX by filling in the dollar amounts for f and v.
  2. What would Quality Tools’ estimated costs be if it had sales of $12,500,000 next month?
  3. What would Quality Tools’ estimated costs be if it had sales of $20,000,000 next month? Why should you feel uncomfortable estimating costs for $20,000,000 in sales?
Answer #1

a)

sales Total cost
Highest activity 16500000 830000
Lowest activity 12000000 650000
change 4500000 180000

Variable cost per $sales = change in cost /change in $ sales

                      = 180000/ 4500000

                        =$ .04 per $ sales

Fixed cost at highest activity=Total cost- variable cost

                   = 830000-(1650000*.04)

                     = 830000- 660000

                     = 170000

Y = 170000+ .04X

B)Estimated cost = 170000+ (.04* 12,500,000)

                         170000+ 500000

                           $ 670000

c)Estimated cost = 170000+ (.04*20,000,000)

                     = 170000+ 800000

                       = 970000

This is so because $ 20,000,000 is a huge amount and not incurred before in any of the reporting period (month)